A particle of mass 3m initially moving with a speed u in the positive x-direction 

collides with a second particle of mass m moving in the opposite direction with an 

unknown speed v. After collision the mass 3m moves along the negative y-direction 

with a speed u/2 and the mass m moves with a speed v in a direction making an angle 

of 45˚ with the positive x-direction. Determine v and v in units of u. Is the collision 

elastic?


1
Expert's answer
2021-01-19T07:11:42-0500

(1)


"3m\cdot\frac{u}{2}=mv'\cdot\sin45°\to v'=\frac{3u}{\sqrt{2}}" . Answer


(2)


"3m\cdot u-mv=mv'\cdot\cos45°"


"3m\cdot u-mv=m\cdot \frac{3u}{\sqrt{2}}\cdot\cos45°\to v=\frac{3u}{2}" . Answer


(3)


before collide


"KE=\frac{3mu^2}{2}+\frac{mv^2}{2}=\frac{3mu^2}{2}+\frac{m(3u/2)^2}{2}=\frac{21}{8}mu^2"


after collide


"KE=\frac{3m\cdot (u/2)^2}{2}+\frac{m\cdot(3u/\sqrt{2})^2}{2}=\frac{21}{8}mu^2"


So, the collision is elastic. Answer